Can you provide answers and solutions for all questions? Thanks in advance !

7 rainbow trout and 2 brook drought are in a pond. If someone fishes with replacement until they catch their first brook trout (then they stop), what is the expected number of fish that they catch in total?
Answer #1

7 rainbow trout

2 brook drought

total 7+2=9

prob(rainbow trout)= 7/9 ==0.78

prob(brook drought)= 2/9  

=0.222

expected 22.2% chanec to catch brook drought

with replacement means any time they have 0.22 chances